SunCar distributes parts for American cars in Belgium. Batteries are a commonly requested item. The lead time to obtain car batteries is two months and the average demand is 100 per month. Lead time demand is normally distributed with a standard deviation of 10 units. Order cost is $75 and holding cost is $10 per year per battery.

The number of orders per year is closest to:

A. 67
B. 5
C. 30
Answer #1

Average monthly demand = 100 units

Annual demand(D) = monthly demand X 12 months = 100 X 12 = 1200 units

Ordering cost(S) = $75

Holding cost(H) = $10

Economic order quantity(EOQ) = Sqrt of (2DS / H)

= Sqrt of [( 2 X 1200 X 75)/10]

= Sqrt of (180000/10)

= Sqrt of 18000

= 134.16 or rounded to 134 units

Number of orders per year = D/EOQ = 1200/134 = 8.96 or rounded to 9